# Congestion-Aware Path Re-routing Strategy for Dense Urban Airspace

### v1

@@ III Congestion Mitigation Methodology
Figure 10: Conflict resolution maneuver executed by UAS with $\psi_{Z}>L$ at time instant $(k-1)\Delta T$ and the descend conditioned maneuver executed by the downstream UAS at time instant $k\Delta T$ on reaching node $\widetilde{Z}$.
The UAS heading in $\alpha_{Z}$ and $\widehat{\delta}_{\mathrm{O}}$ direction would never reach the node $\widetilde{Z}_{\mathrm{UP}}$ simultaneously because of the descend condition and hence, no conflict. However, the UAS transitioning in $\widehat{\delta}_{Z}$ and $\widehat{\gamma}_{Z}$ directions may encounter conflict in some $(k-1)\Delta T$ time instant. The conflict can be resolved by momentarily accelerating and decelerating the UAS. However, once the conflict is resolved, the UAS should reach their next destined cell centers in the $(k+1)\Delta T$ time instant. While UAS is transitioning on the heading-constrained graph, the UAS may encounter zones that are occupied by static obstacles. The smallest rectangular zone cluster that encloses the static obstacle is demarcated as a no-fly zone.
@@
(Presence of static obstacle) For a UAS in zone $Z$, if the $Z_{\mathrm{UP}}$ zone is a no-fly zone, then the UAS perceives $Z_{\mathrm{UP}}$ as always congested. If the $Z_{\mathrm{UD}}$ or $Z_{\mathrm{IN}}$ zone is a no-fly zone, then the UAS perceives $Z_{\mathrm{UD}}$ to be always congested. If the $Z_{\mathrm{ON}}$ zone is a no-fly zone, then the UAS perceives the $Z_{\mathrm{UP}}$ zone as always uncongested (refer Fig. 11).
Figure 11: For UAS in zones marked $Z=1,2,3$ and $4$, corresponding $Z_{\mathrm{UP}}$, $Z_{\mathrm{IN}}$, $Z_{\mathrm{ON}}$ and $Z_{\mathrm{UD}}$ zones are no-fly zones, respectively. Following Rule 7, no UAS in zone $1$ proceeds towards $1_{\mathrm{UP}}$ and no UAS in zone $4$ descends to $4_{\mathrm{UD}}$. Zone $3_{\mathrm{UP}}$ is always perceived as uncongested, hence UAS in $3$ would not move toward $3_{\mathrm{ON}}$. As $2_{\mathrm{IN}}$ is no-fly zone, no UAS in zone $2$ descends to $2_{\mathrm{UD}}$.
As the UAS in the zone marked $1$ in Fig. 11 always perceive $Z_{\mathrm{UP}}$ zone congested, by Rule 2, these UAS would certainly transition along $\gamma_{1}\rightarrow\beta_{1}$ path every successive timeslot. Hence, the $L$-slot look ahead positions of these UAS can be determined and are included in the Definition 2.
@@
In a given timeslot, if the UAS is at Stream 00 node $\widetilde{Z}=(0,Y)$ and finds the $Z_{\mathrm{UP}}=(0,Y+1)$ zone congested, then it randomly opts either of the two $\beta_{0}$ heading directions with probability $\eta$ and $(1-\eta)$ and executes one lateral outward transition in the chosen heading direction.
@@
In a given timeslot, if the UAS present in node $\widetilde{Z}=(\pm 1,Y)$ find the $Z_{\mathrm{UD}}$ zone uncongested. Following Rule 6, either of the UAS in node $(1,Y)$ or $(-1,Y)$ descends to $Z_{\mathrm{UD}}=(0,Y+1)$ with probability $\eta$ and $(1-\eta)$ respectively.
$\eta$ is a UTM design parameter to introduce bias in the traffic distribution about the Stream 00 segment. It is useful when static obstacles or congestion hot spots due to exogenous UAS in the workspace are known apriori.
When every UAS in the finite grid follows the proposed rule-based congestion mitigation strategy, the emerging UAS traffic behavior resembles a parallel air corridor network, with no more than $M$ UAS in any zone and the number of active corridor streams (that is, the UAS traffic spread) adapting to the UAS demand ($\lambda$). When a new source-destination route is being planned for a given $\lambda$, the nominal path for the route, the finite grid, the parameters $S$ (zone dimension), $M$ (minimum number of UAS in the zone for congestion), and $\eta$ are chosen such that the expected traffic spread on this route does not interfere with existing hot spots. The expected UAS traffic spread could be estimated using queuing theoretical models presented in section IV. If $M$ is small, more UAS will transition outward; thus, more traffic spread. $M$ can be increased to lower the spread, compromising intrinsic safety. If $\eta\in[0,1]$ is lower, denser traffic will flow on the left of Stream 00 segment, hence more spread on the left and vice versa. However, if any nominal segment on the route need must pass through the hot spot intersecting the exogenous UAS traffic streams, we assume,
@@
The nominal segment can be intersected by at most one exogenous traffic stream and must intersect orthogonally. The start cell, the end cell, and the finite grid for the nominal segment must be chosen such that the exogenous stream does not pass through any $\gamma\rightarrow\beta$ path in the grid.
The exogenous stream could be in the workspace or out of the plane as depicted by the air corridor in Fig. 3(b).
@@
The traffic flow on the new source-destination route is affected by UAS in the exogenous traffic stream. However, the exogenous stream remains unaffected by the new incoming traffic flow.
The proposed mitigation strategy works even if the above assumptions are violated, as long as the $L$ slot look-ahead positions of UAS in the exogenous streams can be predicted or communicated (refer Fig. 12). Assumptions 1-3 have been stated to develop tractable queuing models for estimating expected traffic spread. Queuing analysis for the more general case is beyond the scope of this paper.
Figure 12: The UAS transitioning on Source2-Destination2 and Source3-Destination3 routes are exogenous traffic streams to UAS transitioning on the Source1-Destination1 route. For $M\leq 4$, from the perspective of a UAS present at node $\widetilde{Z}$, the zone $Z_{\mathrm{UP}}$ is congested. Note that the inter-separation distance between any two UAS present in the square zone is less than $S$ time slots.
The UAS transitioning on the new route does encounter conflicts with UAS in the exogenous streams. However, for these UAS, the proposed congestion mitigation strategy guarantees no more than $M$ conflicts in any zone in the worst case. Here again, the conflicts are resolved by momentarily accelerating and decelerating the UAS, and once the conflict is resolved, the UAS must reach the destined cell center in the same time slot. If the exogenous stream is out of the plane, then in any timeslot at most one exogenous UAS exists in the zone. However, if the exogenous stream is in the workspace, the stream perpendicularly intersects the zone (Source2-Destination2 shown in Fig. 12), thus at most $S$ exogenous UAS can exist in any timeslot. If $\lambda_{e}$ is the probability that an exogenous UAS enters a zone in a given timeslot, then the number of exogenous UAS $(N_{e})$ in the zone is binomially distributed, with the expected number of exogenous UAS $(\mathrm{E}[{N}_{e}])$ in the zone given by,
@@
IV Congestion queuing model
We adopt the discrete-time queuing theory [25, 26, 27, 28] for estimating the expected number of UAS present in any zone $Z=(X,Y)$, $X=\{-X_{e},...,0,...,X_{e}\}$, $Y\in\{1,...,Y_{e}\}$. We define the queuing system $\overline{Z}$, its servers, and queues as follows. Further, we will introduce sufficient supplementary variables so that the resulting queuing system state would be a Markov chain. As discussed in Fig. 7, for UAS transitioning in $\alpha_{Z},\widehat{\beta}_{Z},\widehat{\gamma}_{Z}$ and $\widehat{\delta}_{\mathrm{O}}$ paths, respective $L$ slot look-ahead positions are in zone $Z_{\mathrm{UP}}$ and are responsible for congestion. We associate the zone $Z_{\mathrm{UP}}$ with a queuing system $\overline{Z}$, that comprises of cells belonging to $\alpha_{Z},\widehat{\beta}_{Z},\widehat{\gamma}_{Z}$ and $\widehat{\delta}_{\mathrm{O}}$ and $\gamma_{Z}\rightarrow\beta_{Z}$ paths. Following the rule-based congestion mitigation strategy, the UAS that enter this system of cells ($\overline{Z}$) are arrivals in queuing theory terminology. The downstream UAS that enter the system $\overline{Z}$ via node $\widetilde{Z}$ are referred to as nodal arrivals. Along with nodal arrivals, UAS in zones $Z_{\mathrm{IN}}$ and $Z_{\mathrm{ON}}$ enter the system $\overline{Z}$ via a cell that is not $\widetilde{Z}$ node. These UAS are referred to as non-nodal arrivals. The nodal and non-arrivals are denoted with random variables $\{a_{\alpha,k},a_{\delta,k}\}$, and $\{a_{I,k},a_{O,k}\}$, respectively in Fig. 13. Note that $a_{.,k}=1$ if there is an arrival in time slot $k$, and 00 otherwise. Note that both $a_{\alpha,k}$ and $a_{\delta,k}$ are bernoulli distributed and by Rule 6 (descend condition), in a given time slot $k$ both will not be 1 simultaneously. Thus, we merge the nodal arrivals as $a_{k}=a_{\alpha,k}+a_{\delta_{k}},\ a_{k}\in\{0,1\}$.
@@ IV Congestion queuing model
The $S$ paths summarized in Fig. 9 along with the $\widehat{\delta}_{\mathrm{O}}$ path are the $S+1$ parallel servers of the system. The UAS transitioning on these paths are said to be in service. The UAS that has completed $i$ transitions on any of the $S+1$ paths is said to have received $i$ slots of service, and on reaching the $\widetilde{Z}_{\mathrm{UP}}$ node is said to have departed the $\overline{Z}$ queuing system. A UAS entering service needs $S$ transitions on any of the $S+1$ paths to reach $\widetilde{Z}_{\mathrm{UP}}$ node. Thus, irrespective of the parallel server and the time slot in which the UAS enters service, the service time is $S$ slots (that is $S\Delta T$ time units). Let $w_{i,k}\in\{0,1\},\ i\in\{1,...,S-1\}$ are random variables such that, at time slot $k$, $w_{i,k}=1$ indicates that there exists a UAS that has received $i$ slots of service and $w_{i,k}=0$ otherwise. $w_{1,k}=1$ implies that a UAS has entered service in the previous time slot, and $w_{S-1,k}=1$ implies that a UAS would depart the system $\overline{Z}$ in the next time slot. A UAS that has received $i$ slots of service by time slot $k$ would have received $i+1$ slots of service by $(k+1)$ time slot. Hence,
@@
The UAS in the exogenous stream passing through zone $Z_{\mathrm{UP}}$ also contribute to congestion in $Z_{\mathrm{UP}}$. Corresponding cells through which the stream passes are appended in system $\overline{Z}$. By Assumption 2, the inter-separation distances between UAS present positions, as well as between the $L$ slot look-ahead predicted positions, are geometrically distributed. Statistically, there is no difference between the counting processes related to when the present position and the predicted position enter the zone; hence, they can be interchangeably used for the exogenous stream. The UAS in the exogenous stream on entering zone $Z_{\mathrm{UP}}$ are called exogenous arrivals. According to Assumption 1, the exogenous stream intersects the zone orthogonally. When the exogenous stream is in the workspace, the exogenous arrivals need $S$ transitions to exit the zone (that is, the service time is $S$ slots). Let $e_{i,k}\in\{0,1\},\ i\in\{0,1,...,S-1\}$ be a random variable such that, $e_{0,k}=1$ indicates exogenous arrival at time slot $k$ has entered service. $e_{i,k}=1$ indicates that there exists exogenous arrival that has received $i$ slots of service. When the exogenous stream is out of plane then $i=\{0\}$. If $\lambda_{e}$ is the exogenous arrival rate then,
$e_{i+1,k+1}=e_{i,k}\ \ ,\ \ E_{i+1}(z)=E_{i}(z)$ (6)
$E_{i}(z)=(1-\lambda_{e})+z\ \lambda_{e},\ \ \ i\in\{0,...,S-1\}$ (7)
At any given time slot $k$, the random variables $w_{i,k},\ \ i\in\{1,...,S-1\}$ are independent identically distributed (i.i.d). Note that $w_{1,k+1}$ is dependent on $e_{i,k}\ ,\ i\in\{0,...,S-1\}$ and $w_{1,k}$. But, $w_{1,k}$ is independent of $e_{i,k}$ at time slot $k$. Knowing the PGF of the sum of independent random variables is the product of the PGF of individual random variables; the number of UAS in the system currently in service $u_{k}\in\{0,...,2S-1\}$ and PGF of its limiting probability distribution are given by,
$u_{k}=\sum_{i=1}^{S-1}w_{i,k}+\sum_{i=0}^{S-1}e_{i,k}\ ,\ U(z)=W_{1}(z)^{S-1}E_{0}(z)^{S}$ (8)
$U(z)\triangleq\sum_{n=0}^{2S-1}z^{n}P[u_{k}=n]\ \ ,\ \ k\longrightarrow\infty$ (9)
From the congestion definition 2, $M$ is the minimum number of UAS predicted positions to be present in a zone for the zone to be defined as congested. If in the time slot $k$, $M$ UAS are in service, then the queuing system is said to be busy; otherwise is idle. The queuing system $\overline{Z}$ being busy implies $M$ UAS predicted positions exist in the zone $Z_{\mathrm{UP}}$. Thus,
@@
Let $\theta_{k}$ be a random variable such that $\theta_{k}=0$ indicates the system is busy (that is associated $Z_{\mathrm{UP}}$ zone is congested) and $\theta_{k}=1$ indicates the system is idle (that is associated $Z_{\mathrm{UP}}$ zone is uncongested) at time slot $k$.
$\theta_{k}=\begin{cases}0&\text{if }u_{k}\geq M\\ 1&\text{if }u_{k}<M\end{cases}$ (10)
The PGF of the probability distribution of $\theta_{k}$, as $k\longrightarrow\infty$ is,
$\Theta(z)\triangleq\sum_{n=M}^{2S-1}P[u_{k}=n]+z\sum_{n=0}^{M-1}P[u_{k}=n]$ (11)
From Eqns. 9 and 11, the steady-state probability $\Theta_{0}$ that the zone $Z_{\mathrm{UP}}$ will be congested in time slot $k$ is given by,
$\Theta_{0}$ $=P[\theta_{k}=0]=1-P[\theta_{k}=1]=1-\Theta^{(1)}(0)$
$=1-\sum_{n=0}^{M-1}P[u_{k}=n]=1-\sum_{n=0}^{M-1}\cfrac{U^{(n)}(0)}{n!}$ (12)
where $(n)$ is the $n^{th}$ derivative with respect to $z$. Using Eqn. 8 in Eqn. 12 we get congestion probability in terms of $W_{1}(z)$.
$\Theta_{0}$ $=1-W_{1}(0)^{S-1}E_{0}(0)^{S}$
$\ \ \ -\sum_{n=1}^{M-1}{S-1\choose n}(1-W_{1}(0))^{n}W_{1}(0)^{S-n-1}E_{0}(0)^{S}$
@@
When there are no exogenous arrivals, $E_{0}(0)=1$ in Eqn. 13. $W_{1}(0)=P[w_{i,k}=0]$ is the steady state probability that a UAS arrival will not enter service in a given time slot, and $(1-W_{1}(0))$ is the steady state probability that the arrival would enter service. Eqn. 13 is a forward relation that captures UAS arrivals entering the service are responsible for congestion.
@@
If the $\overline{Z}$ system is busy in time slot $k$, the system remains busy until at least one of the $M$ UAS departs service. The duration (in time slots) for which the system remains busy is referred to as the congestion period. Following Rule 2, the UAS arrivals during the congestion period would successively transition in $\gamma_{Z}\rightarrow\beta_{Z}$ paths. These UAS in the $\gamma_{Z}$ or $\beta_{Z}$ paths are said to be queuing to enter service, with the number of transitions executed before entering service (in timeslots) being the waiting time. The motive is to find a feedback relation that captures congestion being responsible for an arrival entering the queue. It is evident from Fig. 13 that queues of Stream 00 system behave differently from those of Stream $X$ queuing systems. We present queuing models for Stream 00 and Stream $X\neq 0$ systems as follows, highlighting respective differences.
IV-A Stream 00 queuing system

### v2

@@ 3 Congestion Mitigation Methodology
Figure 7: Conflict resolution maneuver executed by the UAS at time instant $k\Delta T$ on reaching node $\widetilde{Z}$.
While UAS is transitioning on the heading-reference graph, the UAS may encounter zones that are occupied by static obstacles. The smallest rectangular zone cluster that encloses the static obstacle is demarcated as a no-fly zone.
@@
(Presence of static obstacle) For a UAS in zone $Z$, if the $Z_{\mathrm{UP}}$ is a no-fly zone, then the $Z_{\mathrm{UP}}$ is treated as a congested zone. If the $Z_{\mathrm{ID}}$ or $Z_{\mathrm{IN}}$ zone is a no-fly zone, then the $Z_{\mathrm{ID}}$ is treated as congested zone. If the $Z_{\mathrm{ON}}$ zone is a no-fly zone, then irrespective of congestion in the $Z_{\mathrm{UP}}$ zone, $Z_{\mathrm{UP}}$ is always considered uncongested so that no UAS performs outwards transitions and heads towards $Z_{\mathrm{ON}}$ (refer Fig. 8).
Figure 8: For UAS in zones marked $Z=1,2,3$ and $4$, corresponding $1_{\mathrm{UP}}$, $2_{\mathrm{ID}}$, $3_{\mathrm{IN}}$, and $4_{\mathrm{ON}}$ zones are no-fly zones, respectively. Following Rule 2, 5 and 7, no UAS in zone $1$ proceeds towards $1_{\mathrm{UP}}$. No UAS in zone $2,3$ descends to $2_{\mathrm{ID}},3_{\mathrm{ID}}$, respectively. Zone $4_{\mathrm{UP}}$ is always perceived uncongested, hence UAS in zone $4$ would never progress towards $4_{\mathrm{ON}}$.
Consider the UAS in zone $Z$ whose $Z_{\mathrm{UP}}$ is a no-fly zone (zone marked $1$ in Fig. 8). As the $Z_{\mathrm{UP}}$ is always perceived congested, these UAS on reaching node $\widetilde{Z}$ would certainly transition along $\beta_{Z}\rightarrow\gamma_{Z}$ path every successive timeslot. Hence, the $L$-slot look ahead positions of these UAS can be determined and are included in the Definition 2.
@@
When UAS in the nominal zone at node $\widetilde{Z}=(0,Y)$ finds the $Z_{\mathrm{UP}}$ zone congested, then the UAS randomly opts either of the two $\gamma_{0}$ paths and executes one lateral transition in the chosen outward direction. The UAS opts for the right outward direction with probability $\eta$ and the left outward direction with probability $(1-\eta)$.
Here, $\eta\in[0,1]$ is a UTM design parameter to introduce bias in the traffic distribution about the nominal segment. It is useful when static obstacles or congestion regions due to exogenous UAS presence in the workspace are known a priori.
@@
Consider UAS transitioning in the Stream(±$(\pm 1)$ zones (one in each zone). If the UAS simultaneously reach respective nodes $\widetilde{Z}=(\pm 1,Y)$ and find the descend condition satisfied, then the probability that UAS in node $(1,Y)$ descends onto the nominal segment, and UAS in node $(-1,Y)$ does not descend is $\eta$. The probability that UAS in node $(-1,Y)$ descends, and UAS in node $(1,Y)$ does not descend is $(1-\eta)$.
When every UAS in the finite grid follows the proposed rule-based congestion mitigation strategy, the emerging UAS traffic behavior resembles a parallel air corridor network, with no more than $M$ UAS in any zone and the number of active corridor streams (that is, the UAS traffic spread) adapting to the UAS demand ($\lambda$). When a new source-destination route is being planned for a given $\lambda$, the nominal path for the route, the finite grid, the parameters $S$ (zone dimension), $M$ (minimum number of UAS in the zone for congestion), and $\eta$ are chosen such that the expected traffic spread on this route does not interfere with existing congested regions. The expected UAS traffic spread could be estimated using queuing theoretical models presented in section 4. If $M$ is small, more UAS will transition outward, leading to more traffic spread. However, $M$ can be increased to lower the spread at the cost of compromising intrinsic safety. If $\eta\in[0.5,1]$, then the probability that UAS opts right outward direction is higher. More traffic will flow to the right of the nominal segment, and hence, there will be more traffic spread to the right. Similarly, if $\eta\in[0,0.5]$, the probability $(1-\eta)$ that UAS opts left outward direction is higher. Hence there would be more traffic spread on the left of the nominal segment.
For each source$(\mathsf{S})$-destination$(\mathsf{D})$ pair existing in the workspace, we have UAS traffic streams emanating at the source and terminating at the destination. The UAS traffic streams between existing $\mathsf{S}$-$\mathsf{D}$ pairs are exogenous traffic streams to the UAS on the new $\mathsf{S}$-$\mathsf{D}$ route. It is preferable that while planning the nominal path for the new route, the nominal path does not intersect exogenous traffic streams. However, in scenarios where a significant reduction in flight time is possible due to shorter path length if the nominal path were passing through exogenous traffic streams (refer Fig. 9a-9c), we recommend that path planning impose the following assumptions.
@@
Each nominal segment in the nominal path can be intersected by at most one exogenous traffic stream. The nominal segment and the exogenous stream must intersect orthogonally. The start cell, the end cell, and the finite grid for the nominal segment must be chosen such that the exogenous stream does not pass through any $\beta\rightarrow\gamma$ path.
@@
The traffic flow on the new source-destination route is affected by UAS in the exogenous traffic stream. However, the UAS in the exogenous traffic stream remain unaffected by the new incoming traffic flow.
The exogenous traffic stream could be in the workspace or out of the plane, as depicted by the air corridor in Fig. 2 front view. The proposed mitigation strategy will be effective even if the above assumptions are violated, as long as the $L$ slot look-ahead positions of UAS in the exogenous streams can be predicted or communicated (refer Fig. 9d). Assumptions 1-3 have been stated to develop tractable queuing models for estimating expected traffic spread. Queuing analysis for the more general case is beyond the scope of this paper.
Figure 9: Path planning where the nominal path of a) the new route $\mathsf{S2}$-$\mathsf{D2}$ avoids the exogenous stream $\mathsf{S1}$-$\mathsf{D1}$, b) the new route $\mathsf{S1}$-$\mathsf{D1}$ avoids the exogenous stream $\mathsf{S2}$-$\mathsf{D2}$, c) the new route $\mathsf{S1}$-$\mathsf{D1}$ intersects the exogenous stream $\mathsf{S2}$-$\mathsf{D2}$. d) The UAS transitioning on $\mathsf{S2}$-$\mathsf{D2}$ and $\mathsf{S3}$-$\mathsf{D3}$ routes are exogenous traffic streams to UAS transitioning on the $\mathsf{S1}$-$\mathsf{D1}$ route. For $M\leq 4$, from the perspective of a UAS present at node $\widetilde{Z}$, the zone $Z_{\mathrm{UP}}$ is congested because of UAS 1,2,3 and 4. Note that the inter-separation distance between any two UAS present in the square zone is less than $S$ timeslots.
The UAS transitioning on the new route does encounter conflicts with UAS in the exogenous traffic streams. However, for these UAS, the proposed congestion mitigation strategy guarantees no more than $M$ conflicts in any zone in the worst case. Here again, the conflicts are resolved by momentarily accelerating and decelerating the UAS, and once the conflict is resolved, the UAS must reach the destined cell center in the same timeslot. If the exogenous traffic stream is out of the plane, then in any timeslot, at most one exogenous UAS exists in the zone. However, if the exogenous stream is in the workspace, the stream intersects the zone orthogonally ($\mathsf{S2}$-$\mathsf{D2}$ shown in Fig. 9d), thus at most $S$ exogenous UAS can exist in any timeslot. If $\lambda_{e}$ is the probability that an exogenous UAS enters a zone in a given timeslot, then the number of exogenous UAS $(N_{e})$ in the zone is binomially distributed, with the expected number of exogenous UAS $(\mathrm{E}[{N}_{e}])$ in the zone given by,
@@
4 Congestion queuing model
We adopt the discrete-time queuing theory (Atencia and Pechinkin (2013); Gao et al. (2004); Wittevrongel and Bruneel (1999); Bruneel and Wuyts (1994)) for estimating the expected number of UAS present in any zone $Z=(X,Y)$, $X=\{-X_{e},...,0,...,X_{e}\}$, $Y\in\{1,...,Y_{e}\}$. As discussed in Fig. 5(b), for UAS transitioning in $\alpha_{Z},\widehat{\beta}_{Z},\widehat{\gamma}_{Z}$ and $\widehat{\delta}_{\mathrm{O}}$ paths, respective $L$ slot look-ahead positions are in zone $Z_{\mathrm{UP}}$ and are responsible for congestion. We associate the zone $Z_{\mathrm{UP}}$ with a queuing system $\overline{Z}$, that comprises cells belonging to $\alpha_{Z},\widehat{\beta}_{Z},\widehat{\gamma}_{Z},\widehat{\delta}_{\mathrm{O}}$ and $\beta_{Z}\rightarrow\gamma_{Z}$ paths. The upstream and the diagonal paths are servers, and the laterally outward paths are queues. Following the rule-based congestion mitigation strategy, the UAS that enter this system of cells ($\overline{Z}$) are arrivals in queuing theory terminology. The downstream UAS that enter the system $\overline{Z}$ via node $\widetilde{Z}$ are referred to as nodal arrivals. Along with nodal arrivals, UAS in zones $Z_{\mathrm{IN}}$ and $Z_{\mathrm{ON}}$ enter the system $\overline{Z}$ via a cell that is not $\widetilde{Z}$ node. These UAS are referred to as non-nodal arrivals. The nodal and non-arrivals are denoted with Bernoulli distributed random variables $\{a_{\alpha},a_{\delta}\}$, and $\{a_{\mathrm{I}},a_{\mathrm{O}}\}$, respectively, in Fig. 10. Note that $a_{.,k}=1$ if there is an arrival in timeslot $k$, and 00 otherwise. Applying Rule 5 (descend condition) in any timeslot $k$, both $a_{\alpha,k},\ a_{\delta,k}$ would simultaneously never be 1. Thus, we merge the nodal arrivals as $a_{k}=a_{\alpha,k}+a_{\delta_{k}},\ a_{k}\in\{0,1\}$.
@@ 4 Congestion queuing model
The $S$ paths summarized in Fig. 6(b) along with the $\widehat{\delta}_{\mathrm{O}}$ path are the $S+1$ parallel servers of the system. The UAS transitioning on these paths are said to be in service. The UAS that has completed $i$ transitions on any of the $S+1$ paths is said to have received $i$ slots of service, and on reaching the $\widetilde{Z}_{\mathrm{UP}}$ node is said to have departed the $\overline{Z}$ queuing system. A UAS entering service needs $S$ transitions on any of the $S+1$ paths to reach $\widetilde{Z}_{\mathrm{UP}}$ node. Thus, irrespective of the parallel server and the timeslot in which the UAS enters service, the service time is $S$ slots (that is $S\Delta T$ time units). Let $w_{i,k}\in\{0,1\},\ i\in\{1,...,S-1\}$ are random variables such that, at timeslot $k$, $w_{i,k}=1$ indicates that there exists a UAS that has received $i$ slots of service and $w_{i,k}=0$ otherwise. Here, $w_{1,k}=1$ implies that a UAS has entered service in the previous timeslot, and $w_{S-1,k}=1$ implies that a UAS would depart the system $\overline{Z}$ in the next timeslot. A UAS that has received $i$ slots of service by timeslot $k$ would have received $i+1$ slots of service by $(k+1)$ timeslot. Hence,
@@
The UAS in the exogenous traffic stream passing through zone $Z_{\mathrm{UP}}$ also contribute to congestion in $Z_{\mathrm{UP}}$. Corresponding cells through which the stream passes are appended in system $\overline{Z}$. By Assumption 2, the inter-separation distances between exogenous UAS current positions, as well as between the $L$ slot look-ahead predicted positions, are geometrically distributed. Statistically, there is no difference between the counting processes related to when the present position and the predicted position enter the zone; hence, they can be interchangeably used for the exogenous stream. The UAS in the exogenous stream on entering zone $Z_{\mathrm{UP}}$ are called exogenous arrivals. According to Assumption 1, the exogenous stream intersects the zone orthogonally. When the exogenous stream is in the workspace, the exogenous arrivals need $S$ transitions to exit the zone (that is, the service time is $S$ slots). Let $e_{i,k}\in\{0,1\},\ i\in\{0,1,...,S-1\}$ be a random variable such that, $e_{0,k}=1$ indicates that exogenous arrival at timeslot $k$ has entered service. Further, $e_{i,k}=1$ indicates that there exists an exogenous arrival that has received $i$ slots of service. When the exogenous stream is out of the plane, then $i=\{0\}$. If $\lambda_{e}$ is the exogenous arrival rate then,
$e_{i+1,k+1}=e_{i,k}\ \ ,\ \ E_{i+1}(z)=E_{i}(z)=(1-\lambda_{e})+z\ \lambda_{e}\ \ ,\ \ \ \ \ \ i\in\{0,...,S-1\}$ (7)
At any given timeslot $k$, the random variables $w_{i,k},\ \ i\in\{1,...,S-1\}$ are independent identically distributed (i.i.d). Note that $w_{1,k+1}$ is dependent on $e_{i,k}\ ,\ i\in\{0,...,S-1\}$ and $w_{i,k}$. But, $w_{i,k}$ is independent of $e_{i,k}$ at timeslot $k$. Knowing the PGF of the sum of independent random variables is the product of the PGF of individual random variables; the number of UAS in the system currently in service $u_{k}\in\{0,...,2S-1\}$ and PGF of its limiting probability distribution are given by,
$u_{k}=\sum_{i=1}^{S-1}w_{i,k}+\sum_{i=0}^{S-1}e_{i,k}\ \ \ ,\ \ \ U(z)=W_{1}(z)^{S-1}E_{0}(z)^{S}\ \ \ ,\ \ \ U(z)\triangleq\sum_{n=0}^{2S-1}z^{n}P[u_{k}=n]\ \ \ ,\ \ \ k\longrightarrow\infty$ (8)
From the congestion Definition 2, $M$ is the minimum number of UAS predicted positions to be present in a zone for the zone to be defined as congested. If in the timeslot $k$, $M$ UAS are in service, then the queuing system is said to be busy; otherwise it is idle. The queuing system $\overline{Z}$ being busy implies $M$ UAS predicted positions exist in the zone $Z_{\mathrm{UP}}$.
@@
Let $\theta_{k}$ be a random variable such that $\theta_{k}=0$ indicates the system is busy (that is associated $Z_{\mathrm{UP}}$ zone is congested) and $\theta_{k}=1$ indicates the system is idle (that is, associated $Z_{\mathrm{UP}}$ zone is uncongested) at timeslot $k$. The PGF of the probability distribution of $\theta_{k}$, as $k\longrightarrow\infty$ is,
$\theta_{k}=\begin{cases}0&\text{if }u_{k}\geq M\\ 1&\text{if }u_{k}<M\end{cases}\ \ \ \ ,\ \ \ \ \Theta(z)\triangleq\sum_{n=M}^{2S-1}P[u_{k}=n]+z\sum_{n=0}^{M-1}P[u_{k}=n]$ (9)
From Eqns. 8 and 9, the steady-state probability $\Theta_{0}$ that the zone $Z_{\mathrm{UP}}$ will be congested in timeslot $k$ is given by,
[…]
where $(n)$ is the $n^{th}$ derivative with respect to $z$. Using Eqn. 8 in Eqn. 10 we get probability $\Theta_{0}$ in terms of $W_{1}(z)$.
$\Theta_{0}\ =$ $1-W_{1}(0)^{S-1}E_{0}(0)^{S}-\sum_{n=1}^{M-1}{S-1\choose n}(1-W_{1}(0))^{n}W_{1}(0)^{S-n-1}E_{0}(0)^{S}$
@@
When there are no exogenous arrivals, $E_{0}(0)=1$ in Eqn. 11. Here, $W_{1}(0)=P[w_{i,k}=0]$ is the steady-state probability that a UAS arrival will not enter service, and $(1-W_{1}(0))$ is the steady-state probability that the arrival would enter service in a given timeslot. Eqn. 11 is a forward relation that captures UAS arrivals entering the service are responsible for congestion.
@@
If the $\overline{Z}$ system is busy in timeslot $k$, the system remains busy until at least one of the $M$ UAS departs service. The duration (in timeslots) for which the system remains busy is referred to as the congestion period. Following Rule 2, the UAS arrivals during the congestion period would successively transition in the $\beta_{Z}\rightarrow\gamma_{Z}$ path. These UAS in the $\beta_{Z}$ or $\gamma_{Z}$ paths are said to be queuing to enter service, with the number of transitions executed before entering service (in timeslots) being the waiting time. The motive is to find a feedback relation that captures congestion being responsible for an arrival entering the queue. It is evident from Fig. 10 that queues associated with Stream$(0)$ zones behave differently from those of Stream$(X)$ zones. We present queuing models for Stream$(0)$ and Stream$(X\neq 0)$ zones as follows, highlighting respective differences.
4.1 Stream$\{0\}$ queuing system